You are asked to prepare 2.3 L of a HCN/NaCNbuffer that has a pH of 9.63 and an osmotic pressure of 1.30 atm at 298 K.

What masses of HCN and NaCN should you use to prepare the buffer? (Assume complete dissociation of NaCN.
Answer #1

Answer:

Given,

To determine the masses of HCN and NaCN

Consider,

Concentration of buffer = P/RT

Substitute the known values in above expression

= (1.30/(0.0821*298))

= 0.05314 M

No of moles of buffer = 0.05314*2.3

= 0.1222 mol

Now,

pH = pka + log(salt/acid)

9.63 = 9.2 + log(x/(0.1222-x))

By solving we get

x = No of mol of NaCN = 0.0891 mol

No of mol of HCN = 0.1222 - 0.0891

= 0.0331 mol

mass of NaCN = 49*0.0891

mass of NaCN = 4.3659 grams

mass of HCN = 0.0331*27

mass of HCN= 0.8937 grams